OpenStudy (anonymous):

I haveone more question. If the lenght of twolegs of a right triangle are: the square root of 3 and the square root of 6 , then the length of the hypotenuse is: ?

OpenStudy (anonymous):

Have you seen the equation: a^2+b^2=c^2?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know what a and b represent?

OpenStudy (anonymous):

is this the legs

OpenStudy (anonymous):

Yep! So then the c represents the hypotenuse. So do you think you would know how to set up the problem?

OpenStudy (anonymous):

not with square roots I don't

OpenStudy (anonymous):

Well, the square roots (surprisingly) actually makes things a little easier. So we start with our original formula: a^2+b^2=c^2 And since you said a and b are the legs, we can substitute the measurements you gave: \[(\sqrt(3))^2+(\sqrt(6))^2=c^2\]

OpenStudy (anonymous):

So technically if you square a square root it sort of cancels with one another. So \[(\sqrt(3))^2=3\]

OpenStudy (anonymous):

and the same idea for \[(\sqrt(6))^2\] which would equal 6.

OpenStudy (anonymous):

So where would you go from there?

OpenStudy (anonymous):

would I go 3+6=c which would be c=9

OpenStudy (anonymous):

Well technically it would be 3+6=9, so 9=c^2. So then you would take the sqaure root of both sides so you have c=?